### v2

# A Scalable Pipeline for Enabling Non-Verbal Speech Generation and Understanding

Abstract
Non-verbal Vocalizations (NVs), such as laughter and sighs, are vital for conveying emotion and intention in human speech, yet most existing speech systems neglect them, which severely compromises communicative richness and emotional intelligence. Existing methods for NVs acquisition are either costly and unscalable (relying on manual annotation/recording) or unnatural (relying on rule-based synthesis). To address these limitations, we propose a highly scalable automatic annotation framework to label non-verbal phenomena from natural speech, which is low-cost, easily extendable, and inherently diverse and natural. This framework leverages a unified detection model to accurately identify NVs in natural speech and integrates them with transcripts via temporal-semantic alignment method. Using this framework, we created and released NonVerbalSpeech-38K, a diverse, real-world dataset featuring 38,718 samples across 10 NV categories collected from in-the-wild media. Experimental results demonstrate that our dataset provides superior controllability for NVs generation and achieves comparable performance for NVs understanding.
Demoshttps://nonverbalspeech38k.github.io/nonverspeech38k/
Datasetshttps://huggingface.co/datasets/nonverbalspeech/nonverbalspeech38k
1 Introduction
Spoken interaction has become a core interface in human-computer communication, driven by the rapid advancement of automatic speech recognition (ASR), text-to-speech (TTS) synthesis, and large language models. These technologies have enabled systems to generate spoken content that is semantically coherent and impressively fluent, while also accurately understanding semantic input  (Zhang et al. 2023; Xie and Wu 2024; Défossez et al. 2024). However, human communication is far richer than just words. Non-verbal Vocalizations (NVs), such as laughter and sighs, are critical components of human spoken communication, carrying essential emotional, intentional, and social cues beyond lexical content (Truong and Van Leeuwen 2007; Cortes et al. 2021; Cowen et al. 2019; Lima et al. 2014). Despite their importance, most current speech systems lack the ability to perceive or express such NVs, hindering the creation of truly human-like interactions.
To genuinely emulate human communication, a speech system should both recognize and produce NVs. This work focuses on two tasks capturing these capabilities: NV speech generation and understanding (Fig. 1 (a)).
Speech synthesis with NVs can be categorized into two distinct approaches. The first is tag-controlled, which inserts explicit tags (e.g., [laughing], [breath]) into text to control the type and position of NVs (Du et al. 2024; Nari-labs 2025). These methods offer fine-grained control but require annotated training data and still struggle with precise timing and occurrence control. The second is spontaneous-style, which predicts NVs directly from contextual cues (Li et al. 2024). While generally producing more natural results, such models also rely heavily on proprietary annotated data, limiting further research.
Figure 1: (a) The task definition, (b) The proposed frame-level detection model, and (c) The TSA method used in the pipeline (Fig. 2).
Table 1: Comparison of Existing Corpora and NVS.
Dataset Lang. # Utter. Source Method # Types Human?
NonVerbalTTS(Borisov et al. 2025) EN $\sim$6K VoxCeleb, Expresso Manual annotation 10 Y
CapSpeech(Wang et al. 2025) EN $\sim$37K LibriTTS-R Rule-based concatenation / overlay 3 N
SMIIP-NV(Wu et al. 2025b) ZH $\sim$17K - Manual recording 3 Y
MNV-17(Mai et al. 2025) ZH $\sim$3K - Manual recording 17 Y
SynParaSpeech(Bai et al. 2025) ZH $\sim$56K - Rule-based synthesis with manual verification 6 Y
NVSpeech(Liao et al. 2025) ZH $\sim$174K Nonspeech7k,Emilia, Genshin,StarRail Manual labeling followed by model annotation 18 Y
NVS (Ours) EN, ZH $\sim$38K In-the-wild Fully automatic annotation 10 N
  • Unlike existing works, the NVS dataset is built using a fully automated annotation pipeline, requiring no human effort and operating directly on in-the-wild data, which allows for straightforward scaling. “# Types” denotes the number of distinct non-verbal types. “Human?” indicates whether human effort is required.
Beyond speech synthesis, understanding NVs is equally important for building truly human-like speech interaction systems. Traditionally, speech processing systems have focused on single-domain tasks such as ASR or sound event classification independently. However, in real-world scenarios, speech and NV sounds naturally co-occur, driving a shift toward unified models that can handle both simultaneously. Recent speech understanding models  (Tang et al. 2024; Chu et al. 2023, 2024; Wu et al. 2025a; Ding et al. 2025) employ multi-task training strategies to jointly perform ASR, sound event classification, and related speech tasks. Although this unified approach integrates multiple tasks within a single framework, effectively combining these capabilities—particularly transcribing NVs information embedded within speech transcription—remains challenging (Shione et al. 2023b, a). Progress is further constrained by the scarcity of large, well-annotated datasets.
As outlined above, the development of both NV speech synthesis and understanding remains fundamentally hindered by the scarcity of publicly available datasets. To address this, several efforts have attempted to incorporate NVs into speech datasets. However, existing datasets often face significant limitations. Specifically, some works rely on manual annotation or human recording, which severely restricts speaker diversity and the scale of the dataset (Mai et al. 2025; Wu et al. 2025b; Borisov et al. 2025). Other approaches use rule-based data augmentation, which may lead to issues with speech naturalness and timbre inconsistency (Wang et al. 2025; Bai et al. 2025). Recent approaches that first annotate a seed dataset and then train a model for automatic expansion (Liao et al. 2025) still rely heavily on costly human effort for the initial labeling.
To address these limitations, we propose a fully automatic annotation framework (Fig. 2) that annotates NVs from in-the-wild audio without introducing costly human effort. Building on this framework, we introduce NonVerbalSpeech-38K (NVS), a large-scale dataset containing naturally occurring NVs along with NV-aware transcriptions. Collected from diverse in-the-wild sources such as animations, movies, variety shows, and radio dramas, NVS contain 38K samples ($\sim$131 hours) covering 10 NV types, including [coughing] and [laughing], as detailed in Fig. 4. Table 1 highlights the key differences between our dataset and existing resources. The key distinctions lie in three aspects: (1) our dataset supports both English and Chinese, covering a broader linguistic scope; (2) its construction is fully automatic, eliminating costly human annotation; and (3) it is derived directly from in-the-wild audio, allowing the dataset to scale continuously.
We evaluate our dataset by fine-tuning F5-TTS (Chen et al. 2024) for NV synthesis and Qwen2-Audio (Chu et al. 2024) and Whisper-Large-V3 (Radford et al. 2023) for end-to-end transcription of both verbal content and NVs. Results show that our dataset offers better controllability for NV generation and achieves competitive performance in NV speech transcription. Demos are available online [^1] https://nonverbalspeech38k.github.io/nonverspeech38k/.
Our main contributions are as follows:
  • A practical, scalable pipeline with two novel components—a unified NV detection model and a precise TSA module—for automatically and accurately detecting and labeling NV expressions from in-the-wild audio.
  • A large-scale, diverse, well-annotated, real-world dataset covering 10 NV categories for expressive NV TTS and precise NV understanding has been released [^2] https://huggingface.co/datasets/nonverbalspeech/nonverbalspeech38k.
  • Fine-tuning F5-TTS with our NVS (TSA) consistently outperforms the strongest baseline (CLAP score $0.110$ (EN) and $0.100$ (ZH); IMOS $3.22$ (EN) and $0.33$ (ZH)) (Table 2), achieving relative gains of $70\%$, $79\%$, $8\%$, and $11\%$, respectively, showing superior NV controllability.
  • Fine-tuning Qwen2-Audio and Whisper-Large-V3 with our NVS (TSA) reduces TPD (sec. 4.2) by 0.92 and 2.64, respectively, compared with the baseline (Table 4), indicating enhanced NV perception.
Figure 2: The proposed NonVerbalSpeech-38K data processing pipeline. Colors other than black in the waveform indicate non-verbal vocalizations.
2 Methodology
This section presents our NonVerbalSpeech-38K annotation pipeline (Fig. 2). It consists of three main stages: Data Preparation and Preprocessing, NV Segment Detection, and Incorporation of NV Tags into Speech Content. We describe each stage in detail below.
2.1 Data Preparation and Preprocessing
As shown in Fig. 2 (Step 1), we first gather in-the-wild data from various sources, such as movies and cartoons. Using the Emilia-Pipeline (He et al. 2024), we preprocess the data through standardization, source separation, and speaker diarization to extract single-speaker segments. Voice activity detection (VAD) removes silence, followed by transcription with Whisper-Large-V3 (Radford et al. 2023) and Paraformer-zh (Gao et al. 2022). Segments with WER $20\%$ or DNSMOS <1.0<1.0 are filtered to ensure high-quality speech. This process also yields segment- and word-level timestamps for further processing.
2.2 Non-Verbal Segments Detection
In this step, we first train the unified frame-level NV detection model and then apply it within the pipeline.
Model Architecture
Inspired by T-UAED (Jiang et al. 2025), we propose a unified frame-level detection model to localize NV events, as illustrated in Fig. 1 (b). Specifically, we employ a pretrained wav2vec-bert 2.0 model  (Barrault et al. 2023) to extract $L$-layer features $F_{i}\in\mathbb{R}^{B\times T\times D}$. To effectively utilize multi-level representations, we apply a learnable weighted sum:
$F=\sum_{i=1}^{L}\alpha_{i}F_{i},\quad\text{s.t. }\sum\alpha_{i}=1,\ \alpha_{i}\geq 0$ (1)
The fused representation $F$ is then passed through a Transformer encoder, followed by a Transformer decoder with learnable queries $Q\in\mathbb{R}^{B\times N_{q}\times D}$:
$F_{\text{enc}}=\text{Encoder}(F),\quad F_{\text{dec}}=\text{Decoder}(Q,F_{\text{enc}})$ (2)
To obtain frame-level predictions, we compute the similarity between decoder and encoder outputs:
$S=\sigma(F_{\text{dec}}F_{\text{enc}}^{\top}),\quad S\in\mathbb{R}^{B\times N_{q}\times T}$ (3)
Here, $\sigma(\cdot)$ denotes the sigmoid function. Finally, the model is trained using binary cross-entropy loss.
Data Preparation for Model Training
To train the frame-level detection model, we construct augmented samples by concatenating or overlaying NVs and speech segments (Fig. 1 (b)). NV clips are collected from widely used sound event datasets including VocalSound  (Gong et al. 2022) , VGGSound  (Chen et al. 2020) , Nonspeech7k  (Rashid et al. 2023) , AudioSet-Strong  (Hershey et al. 2021) , and ESC-50  (Piczak 2015) . To ensure quality, we first apply energy-based VAD to remove silence and split each NV clip into multiple segments. Segments with low CLAP scores (<< 0.4) or very short durations (<< 0.3s) are discarded. The resulting high-quality segments are used to augment training samples, with their distribution shown in Fig. 3. Speech data is drawn from the GigaSpeech M subset ($\sim$1,000 hrs) (Chen et al. 2021) and SEED-TTS-Eval (Anastassiou et al. 2024) . The GigaSpeech test split and the SEED-TTS-Eval set are used exclusively for evaluation.
Experimental Setup and Results
We set $N_{q}=10$, corresponding to the 10 NV types in Fig. 3. Both encoder and decoder have 6 layers. The model is fine-tuned with LoRA (Hu et al. 2022) for 300 epochs on 8 H100 GPUs, using a batch size of 64 per GPU. The learning rate is set to $1\times 10^{-4}$, with a 5% warm-up followed by cosine annealing. See Appendix for details. After training, the model achieves a frame-level F1-score of around 91% on three test splits (Fig. 3). With a similar number of training samples, [gasp] is relatively difficult to perceive and detect, whereas [snore] and [yawn] are easier to distinguish due to their distinctive acoustic patterns. Similarly, [breath] performs slightly worse than [crying] and [throatclearing]. Overall, the frame-level detection model performs relatively poorly on [gasp] and [yawn] due to limited training samples, but performs well on other NV types. Notably, even when trained solely on English speech, the model performs well on Chinese test data, demonstrating strong language-agnostic capabilities and potential for scalability.
Figure 3: Frame-level detection model test results and distribution of training NV segments.
Table 2: Evaluation results of NV speech generation on English (EN) and Chinese (ZH) test sets.
Model CLAP Score $\uparrow$ WER/CER (%) $\downarrow$ DNSMOS $\uparrow$ SSIM $\uparrow$ QMOS $\uparrow$ IMOS $\uparrow$
EN ZH EN ZH EN ZH EN ZH EN ZH EN ZH
CosyVoice2 $0.087$ $0.071$ $3.941$ $5.012$ 3.236{3.236} 3.280{3.280} $0.514$ $0.671$ $3.58_{(\pm 0.06)}$ $3.54_{(\pm 0.07)}$ $2.12_{(\pm 0.06)}$ $2.31_{(\pm 0.08)}$
Dia $0.182$ $0.097$ $15.687$ $171.882$ $2.749$ $2.528$ $0.431$ $0.319$ $3.47_{(\pm 0.13)}$ - $3.27_{(\pm 0.14)}$ -
F5-TTS - - $1.700$ $2.755$ $3.086$ $3.238$ $0.659$ $0.713$ $3.72_{(\pm 0.09)}$ $3.27_{(\pm 0.10)}$ - -
+ CapSpeech $0.125$ $0.121$ $2.791$ $9.856$ $3.009$ $3.153$ $0.570$ $0.669$ $3.66_{(\pm 0.10)}$ $2.40_{(\pm 0.07)}$ $1.79_{(\pm 0.07)}$ $1.88_{(\pm 0.08)}$
+ NVTTS $0.110$ $0.095$ $\mathbf{2.279}$ $7.042$ $3.099$ $\underline{3.248}$ $0.566$ $0.631$ $\mathbf{3.80}_{(\pm 0.08)}$ $3.28_{(\pm 0.08)}$ $22_{(\pm 0.10)} 3.3$$33_{(\pm 0.08)}$
+ NVSpeech $0.092$ $0.094$ $3.525$ $4.726$ $3.040$ $3.214$ $\mathbf{0.596}$ $\underline{0.683}$ $49_{(\pm 0.10)} 3.3$$32_{(\pm 0.07)}$ $2.83_{(\pm 0.11)}$ $3.36_{(\pm 0.09)}$
+ SMIIP-NV $0.257$ $0.281$ 3.501{3.501} 3.910{3.910} $\mathbf{3.132}$ $\mathbf{3.259}$ $0.497$ $0.604$ $75_{(\pm 0.08)} 3.3$$39_{(\pm 0.08)}$ $3.47_{(\pm 0.13)}$ $3.57_{(\pm 0.11)}$
+ SynParaSpeech $0.077$ $0.057$ $3.605$ $4.605$ $3.084$ $3.192$ $0.589$ $\mathbf{0.685}$ $3.25_{(\pm 0.08)}$ $3.00_{(\pm 0.07)}$ $2.88_{(\pm 0.12)}$ $3.12_{(\pm 0.09)}$
+ MNV-17 $0.087$ $0.100$ $5.457$ $4.670$ $3.102$ $\underline{3.248}$ $\underline{0.590}$ $0.667$ $3.43_{(\pm 0.11)}$ $\underline{3.41}_{(\pm 0.08)}$ $2.79_{(\pm 0.10)}$ $3.22_{(\pm 0.07)}$
[…]
+ NVS(TSA) $\mathbf{0.187}$ $\mathbf{0.179}$ $3.498$ $\mathbf{3.101}$ $3.114$ $3.236$ $0.567$ $0.670$ $\underline{3.76}_{(\pm 0.08)}$ $\mathbf{3.43}_{(\pm 0.09)}$ $\mathbf{3.48}_{(\pm 0.08)}$ $\mathbf{3.71}_{(\pm 0.07)}$
  • The best and second-best values in each column are highlighted in bold and underline, respectively. Note that CosyVoice2, Dia, and F5-TTS are excluded as reference models. Gray entries in CLAP Score and QMOS indicate that only a subset of NV types is supported and are therefore excluded from the ranking, see Table 3 for a more detailed comparison. NVTTS: NonVerbalTTS.
Non-Verbal Region Detection Using the Trained Frame-Level Detection Model
After training the frame-level detection model, we apply it to standardized waveforms using a sliding window strategy, as shown in Fig. 2 (Step 2). Standardized waveforms are used because tools like Silero-VAD often miss NV events at the beginning or end of speech segments. To provide sufficient context for detection, we use a window size of 10s and a hop size of 5s. To improve detection quality, we apply three filters: (1) discard segments shorter than 0.3 s; (2) remove segments with very low energy ($35\,\mathrm{dB}$), such as subtle breathing sounds; and (3) discard segments with a CLAP score below 0.3, since higher scores indicate true NV events. These thresholds are determined based on a simple inspection of the data.
2.3 Incorporating Non-Verbal Tags into Speech Content
After filtering out low-quality candidates, we match each NV segment with the nearest speech region obtained from preprocessing (Fig. 2(Step 3(1))). The pseudocode for this process is provided in the Appendix. It should be noted that NV events generally do not lie entirely within speech regions. Therefore, any NV segment located more than 1 second away from the nearest speech region is discarded.
After this, we obtain high-quality speech segments containing NV events, together with NV labels and their start/end timestamps, as well as speech transcriptions with word-level timestamps. To support NV speech generation and understanding, it is necessary to integrate the textual content with NV tags. As shown in Step 3 of Fig. 2, for a segment transcribed as “His funny face made us laugh,” a laughter sound occurs right after “His funny face.” By aligning their timestamps, we obtain the final transcript: “His funny face [laugh] made us laugh.” The algorithm is provided in the Appendix. However, the timestamp-based ordering (TBO) method assumes accurate word timestamps. In practice, ASR timestamps often contain offsets, causing misalignment between the TBO-generated NV-aware transcription and the speech. The Montreal Forced Aligner was expected to improve time stamp precision, but the resulting alignments show greater temporal drift.
To address this, we propose the Temporal-Semantic Alignment (TSA) method (Fig. 1, (c)), yielding two dataset variants (TBO and TSA). We first cut the speech at NV boundaries, run ASR on each segment, and merge the results with NV tags to obtain timing-accurate NV-aware transcriptions. Since this step may introduce semantic inconsistencies, we further align the NV-aware output with the full-segment ASR transcription using edit-distance alignment, combining accurate NV timings with coherent semantics. The final annotation integrates NV tags with the full-segment ASR output. NV events rarely overlap with semantic content, with <B> and </B> co-occurring in less than 2% of cases in the NVS dataset. Note that TSA does not rely on any word-level time stamps, which is a key advantage over TBO.
3 The NonVerbalSpeech-38K Dataset
Using the proposed NonVerbalSpeech-38K pipeline, we construct a dataset from diverse online speech sources (statistics are provided in the Appendix), covering a wide range of natural speaking styles and rich NVs. The initial release contains $\sim$131 hours of English and Chinese speech, totaling about 38,000 samples. As shown in Fig. 4, Chinese samples dominate due to the composition of the source data. Again, the frame-level detection model trained only on English still generalizes well to Chinese, indicating strong cross-lingual scalability. Frequent NV types such as [sigh], [sniff], [laughing], and [coughing] dominate, reflecting their prevalence in daily speech. Less frequent types (e.g., [crying], [yawn]) are also included. Further information and analysis can be found in the Appendix.
Figure 4: Language and NV label distribution in proposed NVS dataset.
Table 3: CLAP Score Differences (Comparison Systems - F5-TTS Fine-Tuned with NVS (TSA)), Broken Down by Non-Verbal Tag.
Model [laughing] [throat] [sigh] [coughing] [sniff] [breath]
English (EN)
CosyVoice2 0.132-0.132 - 0.149-0.149 0.150-0.150 - 0.056-0.056
Dia 0.050-0.050 0.034-0.034 $\mathbf{0.004}$ $\mathbf{0.062}$ 0.085-0.085 -
F5-TTS - - - - - -
+ CapSpeech 0.132-0.132 - - 0.151-0.151 - 0.015-0.015
+ NVTTS 0.160-0.160 0.003-0.003 0.136-0.136 0.043-0.043 0.043-0.043 0.077-0.077
+ NVSpeech 0.131-0.131 - 0.135-0.135 0.144-0.144 - 0.058-0.058
+ SMIIP-NV 0.072-0.072 - - $\mathbf{0.027}$ - -
+ SynParaSpeech 0.117-0.117 0.116-0.116 0.071-0.071 - - -
+ MNV-17 0.171-0.171 0.084-0.084 0.128-0.128 0.142-0.142 0.047-0.047 0.032-0.032
+ NVS(TBO) 0.076-0.076 0.043-0.043 0.105-0.105 0.099-0.099 0.074-0.074 0.051-0.051
Chinese (ZH)
CosyVoice2 0.134-0.134 - 0.124-0.124 0.152-0.152 - 0.098-0.098
Dia 0.105-0.105 $\mathbf{0.004}$ 0.056-0.056 0.132-0.132 0.150-0.150 -
F5-TTS - - - - - -
+ CapSpeech 0.157-0.157 - - 0.127-0.127 - 0.039-0.039
+ NVTTS 0.146-0.146 0.044-0.044 0.108-0.108 0.058-0.058 0.070-0.070 0.078-0.078
+ NVSpeech 0.140-0.140 - 0.103-0.103 0.118-0.118 - 0.055-0.055
+ SMIIP-NV 0.029-0.029 - - $\mathbf{0.058}$ - -
+ SynParaSpeech 0.120-0.120 0.113-0.113 0.060-0.060 - - -
+ MNV-17 0.180-0.180 0.052-0.052 0.092-0.092 0.105-0.105 0.015-0.015 0.033-0.033
+ NVS(TBO) 0.060-0.060 0.042-0.042 0.101-0.101 0.109-0.109 0.057-0.057 0.051-0.051
  • Negative values indicate that F5-TTS fine-tuned with NVS (TSA) outperforms the corresponding system. [throat] denotes [throatclearing], and “-” indicates unsupported NV types for the current comparison system.
4 Experiments
We evaluate the effectiveness of our dataset through two downstream tasks: NV speech generation and NV speech understanding. The details are as follows.
4.1 Experimental Setup for Non-Verbal Speech Generation
Implementation Details
In the NV speech generation task, we use F5-TTS as the base model. To support NV control, we expand the vocabulary of F5-TTS by adding NV tags such as [breath], [sigh]. We fine-tune F5-TTS on seven datasets (Table 1), including the two variants of our NVS dataset. See Appendix for details.
Baselines
In addition to fine-tuning F5-TTS on various datasets, we include two strong baselines, CosyVoice2 (Du et al. 2024) and Dia (Nari-labs 2025), both supporting NV control and generation.
Evaluation
Due to differences in NV support across models and datasets, we select six common NV phenomena for evaluation (Table 3). To evaluate NV speech generation, we use the prompt speech from the seed-tts-eval framework (Anastassiou et al. 2024). For text inputs, tagged sentences are generated using Gemini (Team et al. 2023) and GPT-4o (Hurst et al. 2024). For each NV tag, we construct 175 sentences in both Chinese and English, yielding 1,050 sentences per language. To evaluate system performance, we use standard metrics including DNSMOS, SSIM, and WER/CER to assess speech quality, speaker similarity, and intelligibility. Additionally, we introduce the CLAP Score to measure the presence of intended NV phenomena in synthesized speech. For subjective evaluation, 20 participants rated the samples on QMOS (for overall speech quality) and IMOS (for accuracy in rendering both verbal content and NV tags). See Appendix for details.
4.2 Experimental Setup for Non-Verbal Speech Understaning
Implementation Details
We use Qwen2-Audio and Whisper-Large-V3 as base models, neither of which supports NV-aware transcription. To enable this, we add special NV tokens and fine-tune the models on CapSpeech, NVSpeech, and both NVS variants. See Appendix for details.
Evaluation
To evaluate the model’s ability to integrate NV tags into transcriptions, we use a subset of NonVerbalTTS as the English test set (754 samples, containing [laughing], [breath], and [coughing]), and a subset of SMIIP-NV and MNV-17 as the Chinese test set (784 samples, containing [laughing], [sigh], and [coughing]). All three datasets are manually constructed, ensuring high annotation quality (Table 1). In addition to standard WER/CER (excluding NV tags) and NV-aware WER/CER (NVWER/NVCER, including them), we use two tag-level metrics further evaluating position accuracy: Tag Position Distance (TPD) and its normalized version Normalized Tag Distance (NTD). TPD and NTD are computed after aligning predicted and reference transcriptions using Levenshtein edit distance, where
$\text{TPD}=\frac{1}{n}\sum_{i=1}^{n}|p_{i}-g_{i}|,\quad\text{NTD}=\frac{\text{TPD}}{L},$ (4)
with $p_{i}$ and $g_{i}$ denoting matched tag positions in the prediction and reference and $L$ the reference length. Lower values indicate better accuracy. See Appendix for details.
5 Experimental Results
5.1 Experimental Results on Non-Verbal TTS
Table 2 summarizes NV speech generation results for English and Chinese. Among baselines, CosyVoice2 excels in speech quality (highest DNSMOS and ZH QMOS) but shows limited NV controllability (low CLAP-Score and IMOS). Dia performs well on EN NV metrics (CLAP Score and IMOS) but suffers from poor stability and high failure rates, and is unsupported in Chinese, leading to the exclusion of its ZH subjective scores. F5-TTS is a strong base model, achieving the highest SSIM and lowest WER/CER. However, the pretrained F5-TTS does not support NV generation.
Fine-tuning on NVS (TSA) achieves superior Chinese performance, with top subjective scores, strong objective metrics, and the highest CLAP Score, indicating competitive intelligibility, naturalness, and NV controllability. CapSpeech and NVTTS underperform in ZH CER due to English-only training data. On English tests, NVS(TSA) also performs well, achieving the highest IMOS and CLAP Score, and second-highest DNSMOS and QMOS. CapSpeech attains a high CLAP Score but lowest DNSMOS and IMOS, likely due to unnatural synthetic non-verbal sounds. Tables 3 present CLAP Score comparisons per NV tag. We observe that NVS (TSA) ranks below existing models in only a few cases, demonstrating comprehensive and stable NV controllability. A similar trend is also observed for IMOS, see the Appendix for more details.
Overall, NVS (TSA) performs well across QMOS, IMOS, CLAP Score, ZH CER, and DNSMOS. Compared with NVS (TBO), it shows significant improvements in CLAP Score and IMOS, mainly because TSA provides more accurate annotations, particularly regarding NV positions, highlighting the importance of precise NV labels for controllable NV generation. Nonetheless, the NVS dataset exhibits some limitations in SSIM and English WER, likely due to its size, which we plan to address by further expanding the dataset in future work.
Table 4: EN Evaluation results for NV speech understanding.
Model WER (%) $\downarrow$ NVWER (%) $\downarrow$ TPD $\downarrow$ NTD $\downarrow$
Qwen2-Audio 33.087
+ CapSpeech 20.065 23.261 4.232 0.178
+ NVS(TBO) 16.901 21.952 6.223 0.291
+ NVS(TSA) 17.035 22.099 3.314 0.159
Whisper-Large-V3 17.542
+ CapSpeech 16.037 19.780 3.380 0.167
+ NVS(TBO) 15.469 21.843 8.490 0.396
+ NVS(TSA) 15.503 19.661 0.740 0.055
Table 5: ZH Evaluation results for NV speech understanding.
Model CER (%) $\downarrow$ NVCER (%) $\downarrow$ TPD $\downarrow$ NTD $\downarrow$
Qwen2-Audio 13.868
+ NVSpeech 4.314 4.994 1.530 0.050
+ NVS(TBO) 6.002 8.467 5.419 0.182
+ NVS(TSA) 4.921 6.093 1.737 0.062
Whisper-Large-V3 6.225
+ NVSpeech 4.841 5.166 0.904 0.030
+ NVS(TBO) 5.864 8.141 4.526 0.170
+ NVS(TSA) 5.087 5.844 1.110 0.038
5.2 Experimental Results on Non-Verbal Speech Understaning
Table 4 and 5 compares NV speech understanding performance. On the ZH test set, NVS (TSA) demonstrates competitive performance despite the smaller dataset size and fully automatic construction method without human effort (Table 1), with all metrics remaining low. On the EN test set, it consistently outperforms CapSpeech, which is constructed via rule-based synthesis and limited by dataset diversity. Notably, TSA shows significant improvements over TBO in TPD and NTD across both ZH and EN settings, further validating the superiority of the TSA method, consistent with Table 2.
6 Discussion
Finally, we highlight the scalability of our pipeline:
  • The unified NV detection model is language-agnostic, can generalize to other sounds by increasing query tokens $N_{q}$, and, with CLAP text embeddings, can also support description-based, fine-grained sound detection.
  • Unlike prior work that relies on human effort, our pipeline is fully automated, suggesting that the dataset can be further scaled—a direction for future work. Unlike rule-based synthesis methods, it captures natural NV sounds from diverse sources, which improves NV modeling.
7 Conclusion
In this work, we design a fully automated NV annotation framework featuring two novel components—a unified detection model and the TSA alignment module—avoiding reliance on costly manual labeling and the unnatural outputs caused by rule-based synthesis in prior work. Using this framework, we directly annotate NV phenomena from in-the-wild speech, creating and publicly releasing the NVS dataset, which contains diverse and natural NV events with precise textual labels. Experimental results demonstrate that this dataset effectively enabling existing models to perform NV generation and recognition, particularly exhibiting superior NV controllability.
Appendix A Experimental Setup and Results
In this section, we detail the experimental setup for the frame-level detection model, non-verbal speech generation, and non-verbal speech understanding.
A.1 Experimental Setup and Results for Frame-level Detection Model
We set $N_{q}$ to 10, corresponding to the 10 non-verbal phenomena, and the number of layers $L$ for the pre-trained wav2vec-bert 2.0 model (Barrault et al. 2023) is 25. Both the encoder and decoder are configured with 6 layers, where the decoder excludes the self-attention module. The $Q$ embeddings are randomly initialized. We adopt LoRA (Hu et al. 2022) to fine-tune all linear layers within the self-attention modules of the wav2vec-bert 2.0 model. The model is trained for 300 epochs with a batch size of 64 per GPU, using 8 NVIDIA H100 GPUs. The learning rate is set to $1\times 10^{-4}$ and scheduled with a 5% warm-up phase followed by cosine annealing decay.
After training, the frame-level F1-Score achieves approximately 91% across all three test splits, namely the GigaSpeech test split (Chen et al. 2021), SEED-TTS-Eval Chinese speech (Anastassiou et al. 2024), and SEED-TTS-Eval English speech (Anastassiou et al. 2024). Note that these test sets are also constructed by randomly concatenating or overlaying non-verbal clips with speech segments.
A.2 Experimental Setup and Results for Non-Verbal Speech Generation
Implementation Details
In the non-verbal speech generation task, we use F5-TTS (Chen et al. 2024) as the base model. It is a fully non-autoregressive speech synthesis model that adopts a flow-matching method combined with a transformer architecture. We use the pretrained F5-TTS model, which was trained on the Chinese and English subsets of the Emilia dataset (He et al. 2024), and further fine-tune it on the non-verbal dataset. We use BigVGAN-v2 (Lee et al. 2022) as the vocoder. BigVGAN-v2 is trained on datasets that include diverse types of audio, such as speech in multiple languages, environmental sounds, and musical instruments. This diversity is important for non-verbal speech generation, as it helps the vocoder better handle a wide range of acoustic patterns. To support non-verbal control, we expand the vocabulary of F5-TTS by adding non-verbal tags such as [laughing], [coughing]. Inspired by CapSpeech (Wang et al. 2025), we initialize the embeddings of the new tokens using representations from the CLAP (Wu et al. 2023) text encoder, which is trained on large-scale audio-text alignment data, including non-verbal audio-text pairs.
To retain the basic zero-shot text-to-speech ability of the pretrained model while allowing it to learn non-verbal speech generation, we carefully adjust the total number of training steps based on the scale of each dataset. Based on empirical observations, for all datasets—including ours—but excluding NVSpeech (Liao et al. 2025), SynParaSpeech (Bai et al. 2025), and CapSpeech (Wang et al. 2025), we train the models for 400 epochs to avoid overfitting and maintain fundamental TTS capability. For CapSpeech, we train the model for 50 epochs to achieve the best performance, as extending the training to 400 epochs significantly degrades intelligibility and results in noticeably unnatural speech. NVSpeech and SynParaSpeech are trained for approximately 60k steps, matching the number of training steps used for our dataset, which ensures a fair comparison while preventing overfitting.
All models are optimized using the AdamW optimizer with a base learning rate of $1\mathrm{e}{-5}$, a 10% warm-up phase, and a cosine annealing learning rate schedule. For the newly added non-verbal tokens, a higher learning rate of $1\mathrm{e}{-4}$ is applied to encourage faster adaptation. Fine-tuning is performed on 8 NVIDIA H100 GPUs, with a batch size of 38,400 frames per GPU.
Evaluation
To evaluate system performance, in addition to standard metrics such as DNSMOS [^3] https://github.com/microsoft/DNS-Challenge, SSIM (speaker similarity) [^4] https://github.com/BytedanceSpeech/seed-tts-eval, and WER/CER (Word Error Rate or Character Error Rate) [^5] https://github.com/BytedanceSpeech/seed-tts-eval for assessing speech quality, speaker similarity, and intelligibility, respectively, we also introduce the CLAP Score (Wu et al. 2023) to evaluate whether the synthesized speech contains the intended non-verbal phenomena. Specifically, we use the template sentence “This is a sound of specific non-verbal sound.” (e.g., for laughing, “This is a sound of laughing.”) as the text input to the CLAP (Wu et al. 2023) text tower, and feed the entire synthesized waveform into the CLAP audio tower. For subjective evaluation, we adopt two metrics. The first is QMOS, which assesses the overall quality of the synthesized speech, focusing on intelligibility, audio quality, and naturalness. The second is IMOS, which evaluates whether the audio follows the input text, including non-verbal tags. For example, given the input “[sigh] The work for today is not even halfway done yet.”, the TTS system is expected not only to correctly synthesize the verbal content but also to emit a sigh sound at the appropriate position as indicated by the [sigh] tag. We collected a total of 20 subjective questionnaires. Detailed scoring criteria are provided in the table 6.
Experimental Results
Table 9 presents IMOS comparisons across NV tags. We observe that NVS (TSA) ranks below existing models in only a few cases and achieves consistent leadership in Chinese, demonstrating comprehensive and stable NV controllability.
A.3 Experimental Setup for Non-Verbal Speech Understaning
Implementation Details
In this task, we use Qwen2-Audio and Whisper-Large-V3 as base models, extending their vocabularies to support non-verbal tags (e.g., [laughing], [breath]). For Qwen2-audio, the fine-tuning is performed using LoRA, applied to the linear layers within the self-attention modules of the language model. During training, only the LoRA parameters and the newly introduced tokens are updated, while all other model parameters are frozen. All models are trained using a learning rate of 7.5e-5, batch size 12 per GPU, with 10% warmup and cosine learning rate decay, using 8 NVIDIA H100 GPUs. For Whisper-Large-V3, we full-parameters fine-tuneing the whisper decoder while keep the encoder frozen. All models are trained using a learning rate of 7.5e-5, batch size 24 per GPU, with 10% warmup and cosine learning rate decay, using 8 NVIDIA H100 GPUs.
In the English setting, for a fair comparison, we use only the subset of NonVerbalSpeech-38K containing the three non-verbal tags supported by CapSpeech (i.e., [laughing], [breath], and [coughing]) for training. Details are shown in Table 7. Both datasets (including the two variants of our NonVerbalSpeech-38K) are trained for 5 epochs. In the Chinese setting, we use the overlapping tags in NVSpeech and our dataset (i.e., [laughing], [coughing], [sigh]) for training. Details are shown in Table 8. Both datasets (including the two variants of our NonVerbalSpeech-38K) are trained for 5 epochs. Note that both CapSpeech and NVSpeech have more total samples than our NVS dataset. Even under this setting, our dataset outperforms CapSpeech and achieves performance comparable to NVSpeech, demonstrating its efficiency and effectiveness.
Evaluation
In the English setting, we use a subset of NonVerbalTTS for evaluation, as shown in Table 7. In the Chinese setting, a subset of SMIIP-NV and a subset of MNV-17 are combined and used for evaluation, as shown in Table 8. To simplify the evaluation, we include only samples that contain a single non-verbal tag.
For the evaluation metrics, in addition to standard WER/CER (excluding non-verbal tags) and NV-aware WER/CER (NVWER/NVCER, including them), we introduce two tag-level metrics to further evaluate position accuracy: Tag Position Distance (TPD) and its normalized variant, Normalized Tag Distance (NTD).
  • Tag Position Distance (TPD): To assess the positional accuracy of predicted non-verbal tags, we first align the predicted transcription $T^{(p)}$ with the reference transcription $T^{(r)}$ using the Levenshtein edit distance algorithm. This produces aligned sequences $T_{\text{align}}^{(p)}$ and $T_{\text{align}}^{(r)}$, an example is shown in Table 10. Let $P=\{p_{1},p_{2},\dots,p_{n}\}$ and $G=\{g_{1},g_{2},\dots,g_{n}\}$ denote the positions of the $n$ aligned tags in the predicted and reference sequences, respectively. The TPD is defined as:
    $\text{TPD}=\frac{1}{n}\sum_{i=1}^{n}|p_{i}-g_{i}|.$ (5)
    A lower TPD indicates better alignment of non-verbal tags. In our experiments, both the training and evaluation datasets contain only one tag per utterance to simplify the evaluation and more accurately reflect model performance.
  • Normalized Tag Distance (NTD): To reduce the influence of sequence length, TPD is normalized by the length of the aligned reference transcription:
    $\text{NTD}=\frac{\text{TPD}}{L},\quad\text{where }L=|T_{\text{align}}^{(r)}|.$ (6)
Table 9: IMOS Differences (Comparison Systems - F5-TTS Fine-Tuned with NVS (TSA)), Broken Down by Non-Verbal Tag.
Model [laughing] [throat] [sigh] [coughing] [sniff] [breath]
English (EN)
CosyVoice2 1.40-1.40 - 1.08-1.08 1.60-1.60 - 1.15-1.15
Dia $\mathbf{0.22}$ $\mathbf{0.53}$ 0.95-0.95 1.37-1.37 0.40-0.40 -
F5-TTS - - - - - -
+ CapSpeech 1.81-1.81 - - 2.03-2.03 - 1.38-1.38
+ NVTTS 0.68-0.68 $\mathbf{0.33}$ 0.90-0.90 $\mathbf{0.03}$  + $\mathbf{0.04}$ 0.43-0.43
+ NVSpeech 0.38-0.38 - 0.60-0.60 0.43-0.43 - 0.90-0.90
+ SMIIP-NV 0.11-0.11 - - 0.25-0.25 - -
+ SynParaSpeech 0.46-0.46 0.62-0.62 0.30-0.30 - - -
+ MNV-17 1.35-1.35 0.10-0.10 0.10-0.10 0.13-0.13 0.96-0.96 1.22-1.22
+ NVS(TBO) $\mathbf{0.05}$ 0.17-0.17 0.25-0.25 0.75-0.75 0.38-0.38 0.07-0.07
Chinese (ZH)
CosyVoice2 0.77-0.77 - 1.62-1.62 1.58-1.58 - 1.78-1.78
F5-TTS - - - - - -
+ CapSpeech 1.50-1.50 - - 2.20-2.20 - 1.80-1.80
+ NVTTS 0.50-0.50 0.68-0.68 0.57-0.57 0.03-0.03 0.34-0.34 0.05-0.05  +
+ NVSpeech 0.47-0.47 - 0.10-0.10 0.58-0.58 - 0.02-0.02
+ SMIIP-NV 0.18-0.18 - - 0.06-0.06 - -
+ SynParaSpeech 0.13-0.13 0.60-0.60 0.67-0.67 - - -
+ MNV-17 0.83-0.83 0.26-0.26 0.60-0.60 0.53-0.53 0.32-0.32 0.25-0.25
+ NVS(TBO) 0.03-0.03 0.20-0.20 0.62-0.62 0.05-0.05 0.40-0.40 0.25-0.25
  • Negative values indicate that F5-TTS fine-tuned with NVS (TSA) outperforms the corresponding system. Due to its high CER in Chinese, Dia is excluded from the IMOS ZH evaluation. [throat] denotes [throatclearing], and “-” indicates unsupported NV types for the current comparison system.
Table 10: Example alignment between reference (REF) and model hypothesis (HYP) with non-verbal tag [coughing].
Aligned by Edit Distance
REF [coughing] i just did what **** everybody else does you know after ***** school they kind of look for the next thing and the next thing was you know a levels and
HYP [coughing] i just did what what everybody else does you know after after school he kind of looked for the next thing and the next thing was you know a levels and
  • Insertions (e.g., repeated what, after) and substitutions (e.g., he vs. they) are observed.
Appendix B NonVerbalSpeech-38K Pipeline
B.1 Details on Incorporating Non-Verbal Tags into Speech Content
In this step, each non-verbal segment is matched with the nearest speech region obtained during preprocessing. The pseudocode for identifying the nearest speech region is provided in Algorithm 1.
Algorithm 1 Find Nearest Speech Region for a Non-Verbal Segment
0: Speech regions $S=\{s_{i}\}$; non-verbal segment $n$
0: Nearest $s\in S$ w.r.t. $n$ or discard $n$
[…]
2:for $s\in S$ do
3:  $d\leftarrow\mathrm{distance}(n,s)$
4:  if $n\cap s\neq\varnothing$ then
5:   $(d_{\min},s^{*})\leftarrow(0,s)$
6:   break
7:  else if $d<d_{\min}$ then
8:   $(d_{\min},s^{*})\leftarrow(d,s)$
9:  end if
10:end for
11:if $d_{\min}>\delta$ then
12:  discard $n$
13:else
14:  assign $s^{*}$ to $n$
15:end if
To support non-verbal speech generation and understanding, it is necessary to integrate the textual content with non-verbal tags. The related algorithm is presented in Algorithm 2, which implements the timestamp-based ordering (TBO) method. It is worth noting that state-of-the-art ASR models—including Whisper-Large-V3 (Radford et al. 2023) and Paraformer-zh (Gao et al. 2022)—offer word-level timestamps. In practice, however, we have observed that these timestamps exhibit varying degrees of misalignment, with Paraformer-zh showing particularly noticeable discrepancies. We try using forced alignment tools, such as Montreal Forced Aligner (McAuliffe et al. 2017), to improve alignment. However, due to insertion, substitution, and deletion errors in the transcriptions, no significant improvement is achieved; in some cases, the results are even worse.
Algorithm 2 Insert Non-Verbal Tag into Transcription
0: Non-verbal segment $n=[n_{s},n_{e}]$; non-verbal label $l$; timestamped word list $W=\{(s_{i},e_{i},w_{i})\}_{i=1}^{N}$
0: Updated word list with inserted non-verbal tags
[…]
5:if $i_{e}-i_{s}>1$ then
6:  # Spans multiple words
7:  return $W[0:i_{s}]+[s\_tag]+W[i_{s}:i_{e}]+[e\_tag]+W[i_{e}+1:N]$
8:else
9:  # Single word
[…]
11:end if
Appendix C NonVerbalSpeech-38K Dataset
Figure 5: Speech duration distribution of the proposed NonVerbalSpeech-38K dataset.
Table 11: Source distribution in the original data and the NVS-38K dataset
Source Radio Dramas Comedy Sketches Cartoon Variety Shows Short Plays
Crawled (hrs) 17,400 7,200 3,600 1,400 1,200
NVS-38K (#) 21,668 7,273 5,019 1,217 809
Source Speeches Documentaries Movies Audiobooks Toy Unboxing
Crawled (hrs) 1,079 600 500 263 9
NVS-38K (#) 158 105 1,090 1,375 4
  • “Crawled (hrs)” shows the duration of the original crawled data; “NVS-38K (#)” shows the final sample counts in the proposed NonVerbalSpeech-38K dataset.
Using the proposed NonVerbalSpeech-38K pipeline, we construct the NonVerbalSpeech-38K dataset from a large colillection of speech data crawled from diverse online sources (Table 11). This diversity ensures coverage of a wide range of natural speaking styles and, importantly, includes rich and varied non-verbal phonemes. Radio dramas, comedy sketches, and cartoons contribute the most data, while movies and audiobooks yield the highest processing conversion rates, suggesting richness in non-verbal content. The resulting durations mostly fall within the range of 3 to 30 seconds (Figure 5), primarily because the Emilia (He et al. 2024) pipeline segments speech clips into chunks of 3–30 seconds. This segmentation operation also explains the peak near 30 seconds. Some segments slightly exceeding 30 seconds arise during Step 3 of the NonVerbalSpeech-38K pipeline (Algorithm 1), where non-verbal segments that do not fall within any speech region but are within a 1-second distance threshold are merged with the nearest speech region to maintain contextual continuity.
Limitations: We observe a language imbalance in the current dataset, because the original source datasets are predominantly in Chinese, rather than due to limitations of our pipeline. In fact, our frame-level detection model is trained exclusively on English data, yet it demonstrates strong generalization performance on the Chinese test set. Moreover, since the majority of samples in our dataset are in Chinese, this further substantiates the robustness and cross-lingual generalization ability of the proposed detection model. Thanks to our fully automatic and scalable pipeline, further expansion of the dataset is feasible, and future work will continue to increase both its size and language coverage, helping to mitigate the current language imbalance.